# Learning Network Dismantling without Handcrafted Inputs

### v1

@@ Introduction
[…]
@@
Let $\mathcal{G}$ denote the universe of all possible networks and $P_{\mathcal{G}}$ be the distribution from which a network (or graph) is drawn: $G_{0}=(V_{0},E_{0})\sim P_{\mathcal{G}}$, where $V_{0}$ is the set of nodes and $E_{0}$ is the set of edges between the nodes. At each step $t=0,\cdots,|V_{0}|-1$, the network dismantling policy $\pi(v_{i}|G_{t})$ observes $G_{t}=(V_{t},E_{t})$ and outputs a distribution over $v_{i}\in V_{t}$, from which a node is drawn $v_{t}\sim\pi(v_{i}|G_{t})$ and removed from $G_{t}$ (along with its incident edges), which we formulate as $G_{t+1}=G_{t}\setminus\{v_{t}\}$. With slight abuse of notation, we simplify $v_{t}\sim\pi(v_{i}|G_{t})$ as $v_{t}=\pi(G_{t})$. The standard objective for the network dismantling problem is the area under the curve (AUC) of the relative size of the network’s Largest Connected Component (LCC) to be minimized over the sequence of node removals, which we use to formulate policy optimization:
@@
[…]
@@
The batch $\mathcal{B}$ is sampled from trajectories generated on networks from the same distribution $P_{\mathcal{G}}$ and by the same policy $\pi(v_{i}|G_{t})$ (as in (2)), so, the maximization in Eq. (5) corresponds to a Monte Carlo approximation of the original objective in (2).
@@
To learn the representation of complex networks that is generalizable across all networks in $\mathcal{G}$, MIND employs a GNN-based RL framework, where both the state-action value function $Q(G_{t},v_{i})$ and the policy $\pi(v_{i}|G_{t})$ are parameterized by encoder-decoder neural networks. The encoder GNNs take the adjacency representation of $G_{t}$ as input and extract node embeddings $z_{i}$, which capture the structural role of each node $v_{i}\in V_{t}$. The decoders then map each $z_{i}$ to a scalar score, i.e., the state-action value of removing $v_{i}$ in $Q(G_{t},v_{i})$ decoder, and the probability of selecting $v_{i}$ for removal in $\pi(v_{i}|G_{t})$ decoder. Since the encoders use a permutation-invariant GNN and the decoders are shared across all nodes, this architecture naturally handles networks of varying sizes and calculations of $Q(G_{t},v_{i})$ and $\pi(v_{i}|G_{t})$ for any $(G_{t},v_{i})$ pair.
@@
where $\left[\mathop{\|}_{h=1}^{H}x^{h}\right]$ is a vector concatenation as $[x^{1}\|\cdots\|x^{H}]$, and $\mathrm{MLP}_{\sigma}^{h},\mathrm{MLP}_{\nu}^{h}:\mathbb{R}^{HF}\to(0,1)$ are head-specific neural networks with $\mathrm{sigmoid}$-squashed outputs. Equation (6) uses attention coefficients $\alpha^{h}$ to selectively aggregate messages from different neighbors, similar to the state-of-the-art GATs (Veličković et al. 2018; Brody, Alon, and Yahav 2022). However, GATs do not learn when $e_{i}^{h}=\mathbf{1}_{F}$ for all $i$ and $h$, since softmax-normalization of $\alpha^{h}$ keeps node embeddings identical over message-passing iterations (as demonstrated in Fig. 5).
@@
The first motivation for MIND-MP is the well-known issue of over-smoothing in node embeddings caused by iterative message-passing (Li, Han, and Wu 2018; Oono and Suzuki 2020). In Appendix A, Theorem 1, we show that the embeddings $e_{i}^{(k)}$ for all nodes $v_{i}\in V_{t}$ tend to converge to the principal eigenvector of the message-passing operator as $k$ increases. MIND-MP retains local structural information from early iterations, thereby preserving the diversity of embeddings.
@@
The top three methods, ranked by overall performance across all networks, are MIND (100.0), GDM (104.13), and EI (107.96). The results demonstrate that although other machine learning baselines take advantage of handcrafted inputs, MIND consistently achieves stronger performance across all domains. This highlights that handcrafted initial embeddings, despite boosting the GNN training, do not inherently yield strong dismantling performance. In contrast, MIND, empowered by our proposed MIND-AM and MIND-MP mechanisms (see the GNN Encoder section), is able to identify structurally vital nodes purely from adjacency representation, resulting in an effective network dismantling policy. In technological networks (Fig. 2d), MIND slightly underperforms EI and GND, but still outperforms all other methods, including machine learning baselines (GDM and FINDER). This can be attributed to the limited reach of GNN message-passing in technological networks with very large diameters (e.g., over 100100100 for gridkit-eupowergrid and gridkit-north_america).
Figure 3: The AUC of the dismantling curve for synthetic networks (ER, CM, and SBM) of varying sizes. The scatter plot compares the dismantling performance of all methods normalized for each network relative to MIND, and the bar plot summarizes the overall performance. The AUCs are averaged over 10 realizations.
Figure 4: (a) Relative LCC size during the dismantling of an ER network with 111 k nodes. We compare the node dismantling sequence derived (by PCA) from a set of heuristics with those generated by (b) GDM and (c) MIND; Spearman rank correlation coefficient $R$ and the regression (solid line) with confidence interval (shaded area) are shown on the plots. The heuristic removal sequence is derived from the principal component of GDM’s input node features.
@@
We evaluate MIND on synthetic networks generated by the widely-adopted protocols in prior studies: (i) ER with average degree $\langle d\rangle=4$, (ii) Configuration Model with $\langle d\rangle=4$ and degree distribution $P(d)\sim d^{-2.5}$, and (iii) Stochastic Block Model with group size 100100100, $p_{\text{intra}}=0.1$, and $p_{\text{inter}}=5/|V|$. From each model, we generate networks of size 111 k, 101010 k, and 100100100 k, and evaluate the average AUC over 101010 realizations. Figure 3 shows scatter plots of the AUC of dismantling for all methods, normalized relative to MIND for each network type, and bar plots comparing the overall performances. Note the testing networks in Fig. 3 differ from those that MIND is trained on, in both sizes and methods of generation.
From the results, we observe that MIND significantly outperforms the baselines, except for Stochastic Block Model with 1 $k$ nodes, where the inherent community structures enable the decycling-based method (MS) to achieve a comparable performance to MIND. Notably, although GDM uses the node degree as an input feature, the learned message-passing functions extract structural information that ultimately leads to worse performance than the simple AD across all synthetic networks. Since GDM is trained on the same synthetic networks, this suggests that it overfits to the specific network training set and loses generalizability to similar structures. This highlights the better generalizability of the RL-based (FINDER and MIND) dismantling policies compared to the supervised learning approach (GDM).
For an ER network where nodes are structurally similar, the simple AD method performs considerably better than GDM (Fig.4a). Following this observation, we investigate whether GDM’s degraded performance may be due to an inherent bias towards its handcrafted input features: degree, neighborhood degree statistics, $k$-coreness, and clustering coefficient. Let $X\in\mathbb{R}^{N\times 4}$ be the corresponding feature matrix for the ER network. We combine the input features by projecting $X$ onto the first principle eigenvector of $\frac{1}{N}X^{\top}X$, and order the nodes accordingly to obtain a heuristic dismantling sequence. Figure 4b shows significant correlation between the dismantling sequence of GDM and that of its input features. In contrast, MIND dismantling has weak to no correlation with the heuristic dismantling sequence (Fig. 4c). This corroborates that MIND gains performance by learning the underlying structural importance beyond the standard node heuristics.
@@
Figure 5: The validation AUC during training (mean±$\pm$±std). MIND is compared against: i) GATv2; ii) GCN; iii) MIND without MIND-AM (the all-to-one attention mechanism); and iv) MIND without MIND-MP (the message-profile over iterations).
@@
To assess the effectiveness of our edge-rewiring strategy for diversifying the training networks, we compare MIND with the same model trained on the same networks, only without rewiring. The results are shown in Fig. 6, with bars depicting the effect of the diverse training set on the performance of MIND (shorter bars correspond to higher improvement) on real networks listed in Fig. 2. For each network, the AUC of dismantling with the diversified (rewired) training set is shown as the percentage of the AUC associated with no rewiring (values below 100 indicate that training on rewired networks has led to a better dismantling policy).
@@ Appendix A Convergence of Node Embeddings
Assume $A_{\alpha}$ and $W$ are both diagonalizable with respective dominant eigenvectors $u_{1}\in\mathbb{R}^{N}$ and $v_{1}\in\mathbb{R}^{F}$, whose corresponding eigenvalues are $\lambda_{1}$ and $\mu_{1}$, such that $|\lambda_{1}|>|\lambda_{n}|$ and $|\mu_{1}|>|\mu_{n}|$ for all $n\neq 1$. Then, for any initial embedding $e^{(0)}$ such that $\mathrm{vec}(e^{(0)})$ has nonzero projection onto $u_{1}\otimes v_{1}$:
@@ Appendix B Estimating Structural Heuristics
In this section, we conceptually demonstrate that the node encoding $z_{i}$ generated by MIND is capable of capturing diverse spectral information, even when node features are initialized as constants. For example, we demonstrate that MIND with two heads can approximate the Fiedler vector. (The eigenvector corresponding to the second smallest eigenvalue of the graph Laplacian, often referred to as the Fiedler vector, gives an embedding of nodes that reveals the optimal cut and thus is extremely relevant to network dismantling and similar problems).
@@
which has all eigenvalues lying in the interval (0,1](0,1]( 0 , 1 ]. Moreover, the principal eigenvector $u_{1}$ of $T$ associated with the eigenvalue $\lambda_{1}=1$ is $u_{1}=D^{1/2}\mathbf{1}_{N}$.
@@
i.e., all eigenvalues of $T$ lie in (0,1](0,1]( 0 , 1 ].
In particular, the largest eigenvalue of $T$ is 111 (corresponding to the smallest eigenvalue of $L$, $\lambda=0$). The eigenvector associated with $\lambda=0$ for the normalized Laplacian is proportional to $D^{1/2}\mathbf{1}_{N}$. Hence, the principal eigenvector of $T$ is $u_{1}=D^{1/2}\mathbf{1}_{N}$, up to normalization. ∎
Next, we show that MIND can estimate the Fiedler vector by learning to approximate specific functions in its heads. This follows from the fact that, message-passing operator $T$ defined in (A6) has eigenvalues within the range (0,1](0,1]( 0 , 1 ]. Consequently, repeated message-passing using $T$ gradually attenuates the components of the node embeddings $e^{(k)}$ that align with non-principal eigenvectors. Among these, the component aligned with the Fiedler vector, which corresponds to the second-largest eigenvalue of $T$, decays more slowly than all others except the principal component. As a result, as $k\rightarrow\infty$, only the principal and Fiedler components approximately remain, such that we can approximate the Fiedler vector by performing repeated message-passing with $T$ and subsequently removing the principal component $u_{1}=D^{1/2}\mathbf{1}_{N}$. In particular, MIND with two heads, each with feature dimension $F=1$, is sufficient for such estimation:
Head 1 estimates the principal component $u_{1}$ of $L$ by capturing the node degrees. In the first round of message-passing, MIND can learn the self-weight $W^{1}_{\sigma}=0$ and the neighbor-weight matrix $W^{1}_{\nu}=1$. This yields $e_{i}^{(1)}=d_{i}$, where $d_{i}$ is the degree of node $v_{i}$. Since the degree information can be preserved across all subsequent layers in Head 1 by learning the attention weights $\alpha_{i}^{h}=1$ and $\alpha_{i,j}^{h}=0$ for all heads $h$ and neighbors $j$, we assume that the principal eigenvector $\tilde{u}_{1}=D^{1/2}\mathbf{1}_{N}$ of $T$ is stored in Head 1.
@@
The initial embedding in Head 2 is set to 111 for every node $v_{i}\in V$; that is, the initial embedding vector of Head 2 containing all node embeddings is $e^{(0)}=\mathbf{1}_{N}$. Let $e^{(k)}\in\mathbb{R}^{N}$ denote the vector of node embeddings after the $k$-th message-passing iteration in Head 2. Since $T$ is diagonalizable, we can express the initial embedding vector $e^{(0)}$ as a linear combination of the eigenvectors $u_{1},\dots,u_{N}$ of the matrix $T$:
@@
where $\sqrt{\cdot}$ is the element-wise square root, and $C$ is a scalar constant. This demonstrates that the Fiedler vector can naturally emerge through message-passing followed by simple post-processing by the MLP in (MIND-MP).

### v2

@@ Abstract
Codehttps://github.com/HaozheTian/MIND-ND
Extended versionhttps://arxiv.org/pdf/2508.00706
@@ Introduction
Network dismantling is the problem of finding the sequence of node removals that most rapidly fragments a network into isolated components (Braunstein et al. 2016; Ren et al. 2019). Finding dismantling solutions is equivalent to the identification of vital components of the network system, and has profound real-world applications, such as breaking criminal organizations by arresting the key members (Ribeiro et al. 2018), stopping epidemics with targeted vaccinations (Kitsak et al. 2010; Cohen, Havlin, and Ben-Avraham 2003), ensuring the resilience of healthcare systems via the key providers (Lo Sardo et al. 2019), and preventing wildfires by securing critical locations (Demange et al. 2025). Figure 1 visualizes network dismantling of a real-world social network (Guo, Zhang, and Yorke-Smith 2016) in action, where strategically removing a mere $7\%$ of nodes effectively breaks it into small components.
@@
Let $\mathcal{G}$ denote the universe of all possible networks and $P_{\mathcal{G}}$ be the distribution from which a network (or graph) is drawn: $G_{0}=(V_{0},E_{0})\sim P_{\mathcal{G}}$, where $V_{0}$ is the set of nodes and $E_{0}$ is the set of edges between the nodes. At each step $t=0,\cdots,|V_{0}|-1$, the network dismantling policy $\pi(v_{i}|G_{t})$ observes $G_{t}=(V_{t},E_{t})$ and outputs a distribution over $v_{i}\in V_{t}$, from which a node is drawn $v_{t}\sim\pi(v_{i}|G_{t})$ and removed from $G_{t}$ (along with its incident edges), which we formulate as $G_{t+1}=G_{t}\setminus\{v_{t}\}$. With slight abuse of notation, we simplify $v_{t}\sim\pi(v_{i}|G_{t})$ as $v_{t}=\pi(G_{t})$. The standard objective for the network dismantling problem is to minimize the area under the curve (AUC) of the relative Largest Connected Component (LCC) size of the network over the sequence of node removals, which we use to formulate policy optimization:
@@
where $\text{LCC}(.)$ returns the relative size of the LCC. The optimization problem in (1) can be rewritten as the sum of rewards:
@@
The batch $\mathcal{B}$ is sampled from trajectories generated on networks from the same distribution $P_{\mathcal{G}}$ and by the same policy $\pi(v_{i}|G_{t})$ (as in (2)), therefore, the maximization in Eq. (5) corresponds to a Monte Carlo approximation of the original objective in (2).
@@
To learn the representation of complex networks that is generalizable across all networks in $\mathcal{G}$, MIND employs a GNN-based RL framework, where both the state-action value function $Q(G_{t},v_{i})$ and the policy $\pi(v_{i}|G_{t})$ are parameterized by encoder-decoder neural networks. The encoder GNNs take the adjacency representation of $G_{t}$ as input and extract node embeddings $z_{i}$, which capture the structural role of each node $v_{i}\in V_{t}$. The decoders then map each $z_{i}$ to a scalar score, i.e., the state-action value of removing $v_{i}$ in the $Q(G_{t},v_{i})$ decoder, and the probability of selecting $v_{i}$ for removal in the $\pi(v_{i}|G_{t})$ decoder. Since the encoders use a permutation-invariant GNN and the decoders are shared across all nodes, this architecture naturally handles networks of varying sizes and calculations of $Q(G_{t},v_{i})$ and $\pi(v_{i}|G_{t})$ for any $(G_{t},v_{i})$ pair.
@@
where $\left[\mathop{\|}_{h=1}^{H}x^{h}\right]$ is a vector concatenation as $x^{1}\|\cdots\|x^{H}]$, and $\mathrm{MLP}_{\sigma}^{h},\mathrm{MLP}_{\nu}^{h}:\mathbb{R}^{HF}\to(0,1)$ are head-specific neural networks with $\mathrm{sigmoid}$-squashed outputs. Equation (6) uses attention coefficients $\alpha^{h}$ to selectively aggregate messages from different neighbors, similar to the state-of-the-art GATs (Veličković et al. 2018; Brody, Alon, and Yahav 2022). However, GATs do not learn when $e_{i}^{h}=\mathbf{1}_{F}$ for all $i$ and $h$, since softmax-normalization of $\alpha^{h}$ keeps node embeddings identical over message-passing iterations (as demonstrated in Fig. 5).
@@
The first motivation for MIND-MP is the well-known issue of over-smoothing in node embeddings caused by iterative message-passing (Li, Han, and Wu 2018; Oono and Suzuki 2020). In Appendix A, Theorem 1, we show that the embeddings $e_{i}^{(k)}$ for all nodes $v_{i}\in V_{t}$ tend to converge to the primary eigenvector of the message-passing operator as $k$ increases. MIND-MP retains local structural information from early iterations, thereby preserving the diversity of node embeddings.
@@
The top three methods, ranked by overall performance across all networks, are MIND (100.0), GDM (104.13), and EI (107.96). The results demonstrate that although other machine learning baselines take advantage of handcrafted inputs, MIND consistently achieves stronger performance across all domains. This highlights that handcrafted initial embeddings, despite boosting the GNN training, do not inherently yield strong dismantling performance. In contrast, MIND, empowered by our proposed MIND-AM and MIND-MP mechanisms (see the GNN Encoder section), is able to identify structurally vital nodes purely from adjacency representation, resulting in an effective network dismantling policy. In technological networks (Fig. 2d), MIND slightly underperforms EI and GND, but still outperforms all other methods, including machine learning baselines (GDM and FINDER). This can be attributed to the limited reach of GNN message-passing in technological networks with very large diameters (e.g., over $100$ for gridkit-eupowergrid and gridkit-north_america).
@@
Figure 3: Dismantling performance of MIND and the baseline methods on synthetic networks (ER, CM, and SBM) of varying sizes. The scatter plot compares the dismantling performance of all methods normalized for each network relative to MIND, and the bar plot summarizes the overall performance. The AUCs are averaged over 10 realizations.
Figure 4: (a) Relative LCC size during the dismantling of an ER network with 11 k nodes. We compare the node dismantling sequence derived (by PCA) from a set of heuristics with those generated by (b) GDM and (c) MIND; Spearman rank correlation coefficient $R$ and the regression (solid line) with confidence interval (shaded area) are shown on the plots. The heuristic removal sequence is derived from the principal component of GDM’s input node features.
We evaluate MIND on synthetic networks generated by the widely-adopted protocols in prior studies: (i) ER with average degree $\langle d\rangle=4$, (ii) Configuration Model with $\langle d\rangle=4$ and degree distribution $P(d)\sim d^{-2.5}$, and (iii) Stochastic Block Model with group size $100$, $p_{\text{intra}}=0.1$, and $p_{\text{inter}}=5/|V|$. From each model, we generate networks of sizes $1$ k, $10$ k, and $100$ k, and evaluate the average AUC over $10$ realizations. Figure 3 shows scatter plots of the AUC of dismantling for all methods, normalized relative to MIND for each network type, and bar plots comparing the overall performances. Note the testing networks in Fig. 3 differ from those that MIND is trained on, in both sizes and methods of generation.
From the results, we observe that MIND significantly outperforms the baselines, except for Stochastic Block Model with 1 $k$ nodes, where the inherent community structures enable the decycling-based method (MS) to achieve a comparable performance to MIND. Notably, although GDM uses the node degree as an input feature, the learned message-passing functions extract structural information that ultimately leads to worse performance than the simple AD across all synthetic networks. Since GDM is trained on the same types of synthetic networks, this suggests that it overfits to the specific training set and loses generalizability to similar structures. This highlights the better generalizability of the RL-based (FINDER and MIND) dismantling policies compared to the supervised learning approach (GDM).
For an ER network where nodes are structurally similar, the simple AD method performs considerably better than GDM (Fig.4a). Following this observation, we investigate whether GDM’s degraded performance may be due to an inherent bias towards its handcrafted input features: degree, neighborhood degree statistics, $k$-coreness, and clustering coefficient. Let $X\in\mathbb{R}^{N\times 4}$ be the corresponding feature matrix for the ER network. We combine the input features by projecting $X$ onto the principal eigenvector of $\frac{1}{N}X^{\top}X$, and order the nodes accordingly to obtain a heuristic dismantling sequence. Figure 4b shows a significant correlation between the dismantling sequence of GDM and that of its input features. In contrast, MIND dismantling has weak to no correlation with the heuristic dismantling sequence (Fig. 4c). This corroborates that MIND gains performance by learning the underlying structural importance beyond the standard node heuristics.
@@
[…]
@@
To assess the effectiveness of our edge-rewiring strategy for diversifying the training networks, we compare MIND with the same model trained on the same networks, only without rewiring. The results are shown in Fig. 6, with bars depicting the effect of diversifying the training set on the performance of MIND (shorter bars correspond to higher improvement) on real networks listed in Fig. 2. For each network, the AUC of dismantling with the diversified (rewired) training set is shown as the percentage of the AUC associated with no rewiring (values below 100 indicate that training on rewired networks has led to a better dismantling policy).
The results demonstrate that rewiring the training networks yields an overall performance improvement on real networks. To analyze the results, we refer to the assortativity and modularity of the real networks in Table 4 in Appendix E. The most significant performance gains are observed for highly modular networks. For instance, in roads-california (third-to-last bar in the lower panel of Fig. 6), which has a modularity of $0.975$, training with rewired networks led to an over $80\%$ reduction in the AUC of dismantling. This improvement is likely linked to the rewiring-induced modularity in the otherwise non-modular synthetic networks. We also observe notable performance gains for networks with strong disassortativity. For example, in munmun_twitter_social (seventh-to-last bar in the upper panel of Fig.6), which has a degree assortativity of 0.878-0.878, the AUC is reduced by $40\%$. Although the LPA and Copying Model naturally produce slightly disassortative networks, our diversifying rewirings enable the model to interact with a much wider range of (dis)assortative mixings and thereby significantly enhance the learned embedding and policy by increased exposure to different topologies.
@@ Appendix A Convergence of Node Embeddings
Assume $A_{\alpha}$ and $W$ are both diagonalizable with respective primary eigenvectors $u_{1}\in\mathbb{R}^{N}$ and $v_{1}\in\mathbb{R}^{F}$, whose corresponding eigenvalues are $\lambda_{1}$ and $\mu_{1}$, such that $|\lambda_{1}|>|\lambda_{n}|$ and $|\mu_{1}|>|\mu_{n}|$ for all $n\neq 1$. Then, for any initial embedding $e^{(0)}$ such that $\mathrm{vec}(e^{(0)})$ has nonzero projection onto $u_{1}\otimes v_{1}$:
@@ Appendix B Estimating Structural Heuristics
In this section, we conceptually demonstrate that the node encoding $z_{i}$ generated by MIND is capable of capturing diverse spectral information, even when node features are initialized as constants. For example, we demonstrate that MIND with two heads can approximate the Fiedler vector. (The eigenvector corresponding to the second smallest eigenvalue of the graph Laplacian, often referred to as the Fiedler vector, gives an embedding of nodes that reveals the optimal cut and thus is extremely relevant to network dismantling and similar problems.)
@@
which has all eigenvalues lying in the interval $(0,1]$. Moreover, the primary eigenvector $u_{1}$ of $T$ associated with the eigenvalue $\lambda_{1}=1$ is $u_{1}=D^{1/2}\mathbf{1}_{N}$.
@@
i.e., all eigenvalues of $T$ lie in $(0,1]$.
In particular, the largest eigenvalue of $T$ is $1$ (corresponding to the smallest eigenvalue of $L$, $\lambda=0$). The eigenvector associated with $\lambda=0$ for the normalized Laplacian is proportional to $D^{1/2}\mathbf{1}_{N}$. Hence, the primary eigenvector of $T$ is $u_{1}=D^{1/2}\mathbf{1}_{N}$, up to normalization. ∎
Next, we show that MIND can estimate the Fiedler vector by learning to approximate specific functions in its heads. This follows from the fact that, message-passing operator $T$ defined in (A6) has eigenvalues within the range $(0,1]$. Consequently, repeated message-passing using $T$ gradually attenuates the components of the node embeddings $e^{(k)}$ that align with non-primary eigenvectors. Among these, the component aligned with the Fiedler vector, which corresponds to the second-largest eigenvalue of $T$, decays more slowly than all others except the primary eigenvector. As a result, as $k\rightarrow\infty$, only the primary and Fiedler vectors approximately remain, such that we can approximate the Fiedler vector by performing repeated message-passing with $T$ and subsequently removing the primary eigenvector $u_{1}=D^{1/2}\mathbf{1}_{N}$. In particular, MIND with two heads, each with feature dimension $F=1$, is sufficient for such estimation:
Head 1 estimates the primary eigenvector $u_{1}$ of $L$ by capturing the node degrees. In the first round of message-passing, MIND can learn the self-weight $W^{1}_{\sigma}=0$ and the neighbor-weight matrix $W^{1}_{\nu}=1$. This yields $e_{i}^{(1)}=d_{i}$, where $d_{i}$ is the degree of node $v_{i}$. Since the degree information can be preserved across all subsequent layers in Head 1 by learning the attention weights $\alpha_{i}^{h}=1$ and $\alpha_{i,j}^{h}=0$ for all heads $h$ and neighbors $j$, we assume that the primary eigenvector $\tilde{u}_{1}=D^{1/2}\mathbf{1}_{N}$ of $T$ is stored in Head 1.
@@
The initial embedding in Head 2 is set to $1$ for every node $v_{i}\in V$; that is, the initial embedding vector of Head 2 containing all node embeddings is $e^{(0)}=\mathbf{1}_{N}$. Let $e^{(k)}\in\mathbb{R}^{N}$ denote the vector of node embeddings after the $k$-th message-passing iteration in Head 2. Since $T$ is diagonalizable, we can express the initial embedding vector $e^{(0)}$ as a linear combination of the eigenvectors $u_{1},\dots,u_{N}$ of the matrix $T$:
@@
where $C$ is a constant scalar. The inner-product $\tilde{u}_{1}^{\top}e^{(k)}$ aggregates information from all nodes. Within the message-passing framework, this is made possible by our omni-node $v_{o}$, which is an extra node connected to every node in the original network.
Overall, this section demonstrates that, with constant input node features, the Fiedler vector can naturally emerge from message-passing, enabled by the MIND-MP module and the omni-node.